# CrossGen: Learning and Generating Cross Fields for Quad Meshing

### v1

@@ 4.2. Analysis
Robustness to Missing and Sparse Input.
[…]
Figure 12. Quad meshes generated by our method under varying input conditions, where the normal of our input point cloud is obtained using Xu et al. (2023) method. From left to right: (a) the original point cloud with 150K points; (b) a point cloud with missing regions (highlighted in light violet); and (c) a sparse point cloud with 10K points.
A key strength of CrossGen lies in its robustness to diverse point cloud inputs, enabled by the flexibility of using point clouds. Two common scenarios are incomplete and sparsely sampled data. In real-world settings, 3D scans frequently suffer from missing regions or low-resolution sampling. Unlike traditional methods that rely on complete surface meshes, CrossGen remains effective even when substantial portions of the input geometry are absent. By leveraging a locality-aware latent space, our model can infer global structure and cross fields from limited data. As shown in Fig. 12, CrossGen successfully reconstructs coherent surfaces and high-quality quad meshes even when inputs contain missing regions (b) or are reduced to as few as 10K points (c), demonstrating strong resilience and practicality in challenging data conditions.
@@
Robustness to Rotation
Another important feature of CrossGen is its robustness to random rotations. Unlike most existing methods that rely on shapes in a canonical pose, our method can process shapes with random pose inputs. By leveraging the learned local latent space representation, CrossGen can handle arbitrary orientations, making it more flexible and applicable in dynamic scenarios where the pose of the input shapes may vary without a requirement of a predefined orientation, as shown in Fig. 15.
@@ 5. Ablation Study
Model setting Rotation Ours
AE 0.34 0.05
Time (ms) 73.26 66.75

### v2

@@ 4.2. Analysis
Robustness to Missing, Uneven, and Sparse Inputs.
A key strength of CrossGen lies in its robustness to diverse point cloud inputs, enabled by the flexibility of using point clouds. Three common scenarios are incomplete, uneven, and sparsely sampled data. In real-world settings, 3D scans frequently suffer from missing regions, uneven sampling, or low-resolution sampling. Unlike traditional methods that rely on complete surface meshes, CrossGen remains effective even when substantial portions of the input geometry are absent. By leveraging a locality-aware latent space, our model can infer global structure and cross fields from limited data. As illustrated in Fig. 12, CrossGen successfully reconstructs coherent surfaces and high-quality quadrilateral meshes under challenging conditions, including inputs with missing regions (b), uneven sampling (c), and varying degrees of sparsity (d). We note, however, that excessively sparse inputs may result in the loss of detail, leading to overly smooth reconstructions.
@@
Robustness to Rotation
Another important feature of CrossGen is its robustness to random rotations. Data-driven methods, such as Dielen et al. (2021), often require shapes to be in a canonical pose. In contrast, by leveraging a learned local latent space representation and data augmentations, CrossGen can handle arbitrary orientations, making it robust to input pose variations such as rotations and non-rigid deformations, as illustrated in Fig. 15.
@@ 5. Ablation Study
Angular Error (AE) $\downarrow$ Time (ms) $\downarrow$
Rotation 0.34 73.26
Ours 0.05 66.75
6. Limitations
@@ 7. Conclusion
Acknowledgements.
The authors would like to thank the anonymous reviewers for their valuable comments and suggestions. This work was partially supported by the Innovation and Technology Commission of the HKSAR Government under the InnoHK initiative, the Hong Kong RGC (Ref. T45-205/21-N), the National Key R&D Program of China (2022YFB3303200), the National Natural Science Foundation of China (U23A20312, 62272277), the German Research Foundation within the Gottfried Wilhelm Leibniz program, and the National Science Foundation (OAC-1845962, OAC-1910469, and OAC-2311245).